OpenStudy (anonymous):

find the sum of n terms of a series whose nth term is given by n^2+2^n the answer i am getting is 1/6 n(n+1)(2n+1) + 2 [1. {(2^n) -1}/(2-1) ] but in the book it is given as 1/6 n(n+1)(2n+1) + (2^n) -1...which is correct ?

OpenStudy (anonymous):

your answer is correct \(\large \sum \limits_{n=1}^n 2^n = 2(2^n-1)\)

OpenStudy (anonymous):

thanks @beccaboo022a